# Nonlinear Magnetically Charged Black Holes with Phantom Global Monopoles: Thermodynamics, Geodesics, Quasinormal Modes, and Grey-Body Factors

### v2

@@ Abstract
We study the properties of a nonlinear magnetic-charged black hole in the presence of a phantom global monopole. By incorporating nonlinear electrodynamics (NED) and exotic scalar fields, we derive an exact black hole solution and analyze its geometric structure, causal properties, and thermodynamic behavior. We examine how the presence of a phantom global monopole modifies the black hole’s Hawking temperature, entropy, and stability conditions, revealing significant changes in its phase structure. Additionally, we investigate the geodesic motion of test particles. The quasinormal mode (QNM) spectrum is computed using the WKB approximation and Pöschl-Teller potential method, providing insights into the perturbative stability of the system. Furthermore, we analyze the grey-body factors that characterize radiation emission, highlighting their dependence on black hole parameters. Our findings indicate that the interplay between phantom energy, NED, and global monopoles introduces observable deviations in strong-field astrophysical phenomena. These results offer potential signatures for testing modified gravity theories and contribute to a deeper understanding of black hole physics in exotic field environments.
@@ 1 Introduction
Black holes in modified gravity theories and exotic matter environments exhibit profound deviations from general relativity (GR), offering unique opportunities to probe high-energy astrophysics and quantum gravity effects [1, 2, 3, 4, 5]. While classical black hole solutions in GR are well established, alternative theories predict significant modifications in strong-field regimes. These deviations can manifest in astrophysical observations such as black hole shadows, gravitational waves, and high-energy emissions, providing crucial insights into physics beyond GR. Additionally, the presence of exotic fields and nonlinear electrodynamics (NED) around black holes alters their thermodynamic stability, causal structure, and radiation properties. Investigating these effects is essential for distinguishing modified gravity scenarios from standard astrophysical processes. For a comprehensive review of these modifications and their implications, see [3] and references therein.
@@
Among the exotic fields that may alter black hole physics, phantom energy is of particular interest due to its ability to violate the null energy condition, leading to counterintuitive consequences [6]. Unlike conventional dark energy models, phantom fields have an equation of state parameter less than $-1$, which can significantly impact black hole thermodynamics and stability. Observational data have not ruled out phantom energy [11], making it an active research topic across various branches of physics. Notably, phantom energy accretion reduces black hole mass, potentially violating cosmic censorship and the generalized second law of thermodynamics [12, 13, 14]. In Reissner-Nordström (RN)-like black holes, this process increases the charge-to-mass ratio beyond unity, modifying standard black hole dynamics [15, 16]. Furthermore, phantom fields affect gravitational lensing properties [17], strong lensing effects, and light propagation [18, 19, 20]. Beyond gravitational effects, phantom fields alter black hole thermodynamics by increasing temperature as mass decreases, influencing the evaporation process via Hawking radiation [21]. However, the stability of phantom-dominated black holes remains uncertain, as negative-energy fluxes can introduce instabilities in the event horizon structure [22, 25, 23, 24, 26]. Further studies in various black hole models, including BTZ black holes in commutative [25] and noncommutative [26] spaces, have highlighted the need for a deeper investigation into their thermodynamic fate and role in an accelerating universe.
@@
Another fundamental extension of GR involves NED, which modifies Maxwell’s theory to account for strong-field effects [30, 29]. Such modifications arise naturally in high-energy regimes, particularly around magnetized black holes or intense electromagnetic fields [31, 32]. Several studies indicate that NED can lead to regular black hole solutions, mitigating singularities while preserving essential thermodynamic properties [35, 34, 33]. The thermodynamics of black holes within NED frameworks have been extensively analyzed, revealing rich phase structures, critical behavior, and modified heat capacities [36, 37, 39, 38]. However, when combined with exotic fields such as phantom energy, NED introduces additional complexities affecting black hole stability, causal structure, and accretion dynamics [41, 40]. Recent research has explored magnetically charged black holes in NED coupled to phantom fields, uncovering novel regular solutions and black bounce geometries [42]. Additionally, investigations into rotating and nonlinearly charged black holes with anisotropic matter fields have demonstrated significant deviations from standard GR solutions [43]. These modifications influence observable phenomena, including black hole shadows, gravitational lensing, and quasinormal modes (QNMs), providing potential avenues for testing deviations from classical GR [44]. Lower-dimensional studies, such as those on BTZ black holes, further highlight the unique thermodynamic and geometric properties introduced by NED [39]. As research continues to explore the interplay between NED and unconventional matter fields, these studies offer valuable insights into quantum gravity corrections, alternative compact objects, and broader modified gravity theories.
@@
Despite extensive research on phantom energy, nonlinear electrodynamics, and global monopoles individually, their combined impact on black hole physics remains largely unexplored. This study aims to fill this gap by systematically investigating how their interplay alters black hole thermodynamics, stability, and observational features. This paper is structured as follows: Section 2 presents the derivation of the black hole solution and its geometric properties, while Section 3 analyzes its thermodynamic behavior. Section 4 investigates the timelike geodesics and the motion of test particles in the black hole spacetime. Section 5 examines QNMs using the sixth-order WKB approximation, cross-checked with the Pöschl-Teller potential. Finally, Section 6 analyzes grey-body factors and their influence on radiation emission, before we conclude in Section 7 with a summary and future outlook.
2 Spacetime Geometry and Black Mole Metric
@@
Following the approach in [45], we assume the asymptotic behavior $h(r)\to 1$. Under this assumption, the energy-momentum tensor associated with the global monopole takes the form:
@@
On the other hand, when coupling Einstein gravity with a nonlinear electromagnetic field in the presence of a phantom global monopole, the field equations governing the system take the form:
@@
where the nonlinear electrodynamics (NED) Lagrangian density is given by:
Here, $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$, while $m$ and $Q$ correspond to the mass and magnetic charge of the system, respectively.
@@
Using the previously derived equations, the time component of Eq. (6) simplifies to:
$-\frac{2}{r^{2}}\frac{dM\left(r\right)}{dr}=-\frac{6mQ^{3}}{\left(r^{3}+Q^{3}\right)^{2}}+8\pi\frac{\xi\eta^{2}}{r^{2}}.$ (13)
Integrating this expression, we obtain the metric function that describes a black hole with a nonlinear electromagnetic field in the presence of a phantom global monopole:
@@
In the special case where $\xi\rightarrow 0$, this solution reduces to a black hole with a nonlinear magnetic charge. On the other hand, setting $Q\rightarrow 0$ retrieves the Schwarzschild black hole solution with a phantom global monopole [49]. To investigate the curvature behavior of this spacetime, we compute the Ricci scalar $\mathbf{R}$ and the Kretschmann scalar $\mathbf{K}$, given by:
@@
Both the Ricci scalar $\mathbf{R}$ and the Kretschmann scalar $\mathbf{K}$ exhibit divergences at $r\rightarrow 0$, signifying the presence of a curvature singularity at the origin. However, in the asymptotic limit $r\rightarrow\infty$, both scalars vanish, implying that the spacetime flattens at large distances.
@@
Moreover, the presence of the parameter $\xi$ (implicitly affecting $\eta$) influences the global behavior of the solution, indicating that a sufficiently strong monopole field can significantly alter black hole thermodynamics and causal structure. In particular, the limiting behavior at large $r$ remains consistent, reinforcing the asymptotic flatness of the spacetime. However, near the black hole horizon, the modifications become more pronounced, potentially impacting observational signatures such as gravitational lensing and black hole shadows.
@@ 3 Thermodynamics
where $\kappa$ represents the surface gravity at the event horizon. To evaluate this temperature, we first determine the black hole mass $m$ in terms of the horizon radius $r_{H}$ by solving the condition $f(r_{H})=0$. This leads to the following expression:
@@
$T_{H}=\frac{\left(1-8\pi\eta^{2}\xi\right)}{4\pi r_{H}}\frac{\left(r_{H}^{3}-2Q^{3}\right)}{\left(r_{H}^{3}+Q^{3}\right)}.$ (19)
For the Hawking temperature to remain both real and positive, two conditions must be satisfied. The first constraint stems from the nonlinear electromagnetic field, establishing a lower bound on the event horizon radius $r_{H}$ in terms of the magnetic charge $Q$:
@@
Figure 2: Hawking temperature $T_{H}$ as a function of the horizon radius $r_{H}$ for different values of the symmetry-breaking scale $\eta$. The black hole parameters are set to $Q=0.5$ and $m=1$. The effect of the phantom global monopole ($\xi=-1$) leads to an increase in temperature, whereas the ordinary global monopole ($\xi=+1$) lowers it.
@@
where $A=4\pi r_{H}^{2}$ represents the area of the event horizon. It is noteworthy that the entropy of the black hole remains the same for both the ordinary and phantom cases, suggesting that the presence of the phantom global monopole does not affect the black hole’s entropy.
@@
$G=\frac{\left(1-8\pi\eta^{2}\xi\right)\left(r_{H}^{6}+8Q^{3}r_{H}^{3}-2Q^{6}\right)}{4_{H}r_{H}^{2}\left(r_{H}^{3}+Q^{3}\right)}.$ (27)

### v3

@@ Abstract
We study the properties of a nonlinear magnetic-charged black hole in the presence of a phantom global monopole. By incorporating nonlinear electrodynamics (NLE) and exotic scalar fields, we derive an exact black hole solution and analyze its geometric structure, causal properties, and thermodynamic behavior. We examine how the presence of a phantom global monopole modifies the black hole’s Hawking temperature, entropy, and stability conditions, revealing significant changes in its phase structure. Additionally, we investigate the geodesic motion of test particles. The quasinormal mode (QNM) spectrum is computed using the WKB approximation and Pöschl-Teller potential method, providing insights into the perturbative stability of the system. Furthermore, we analyze the grey-body factors that characterize radiation emission, highlighting their dependence on black hole parameters. Our findings indicate that the interplay between phantom energy, NLE, and global monopoles introduces observable deviations in strong-field astrophysical phenomena. These results offer potential signatures for testing modified gravity theories and contribute to a deeper understanding of black hole physics in exotic field environments.
@@ 1 Introduction
Black holes in modified gravity theories and exotic matter environments exhibit profound deviations from general relativity (GR), offering unique opportunities to probe high-energy astrophysics and quantum gravity effects [1, 2, 3, 4, 5]. While classical black hole solutions in GR are well established, alternative theories predict significant modifications in strong-field regimes. These deviations can manifest in astrophysical observations such as black hole shadows, gravitational waves, and high-energy emissions, providing crucial insights into physics beyond GR. Additionally, the presence of exotic fields and nonlinear electrodynamics (NLE) around black holes alters their thermodynamic stability, causal structure, and radiation properties. Investigating these effects is essential for distinguishing modified gravity scenarios from standard astrophysical processes. For a comprehensive review of these modifications and their implications, see [3] and references therein.
@@
Among the exotic fields that may alter black hole physics, phantom energy is of particular interest due to its ability to violate the null energy condition, leading to counterintuitive consequences [6]. Unlike conventional dark energy models, phantom fields have an equation of state parameter less than 1-1, which can significantly impact black hole thermodynamics and stability. Observational data have not ruled out phantom energy [11], making it an active research topic across various branches of physics. Notably, phantom energy accretion reduces black hole mass, potentially violating cosmic censorship and the generalized second law of thermodynamics [12, 13, 14]. In Reissner-Nordström (RN)-like black holes, this process increases the charge-to-mass ratio beyond unity, modifying standard black hole dynamics [15, 16]. Furthermore, phantom fields affect gravitational lensing properties [17], strong lensing effects, and light propagation [18, 19, 20]. Beyond gravitational effects, phantom fields alter black hole thermodynamics by increasing temperature as mass decreases, influencing the evaporation process via Hawking radiation [21]. However, the stability of phantom-dominated black holes remains uncertain, as negative-energy fluxes can introduce instabilities in the event horizon structure [22, 25, 23, 24, 26]. Further studies in various black hole models, including BTZ black holes in commutative [25] and noncommutative [26] spaces, have highlighted the need for a deeper investigation into their thermodynamic fate and role in an accelerating universe.
@@
Another fundamental extension of GR involves NLE, which modifies Maxwell’s theory to account for strong-field effects [30, 29]. Such modifications arise naturally in high-energy regimes, particularly around magnetized black holes or intense electromagnetic fields [31, 32]. Several studies indicate that NLE can lead to regular black hole solutions, mitigating singularities while preserving essential thermodynamic properties [35, 34, 33]. The thermodynamics of black holes within NLE frameworks have been extensively analyzed, revealing rich phase structures, critical behavior, and modified heat capacities [36, 37, 39, 38]. However, when combined with exotic fields such as phantom energy, NLE introduces additional complexities affecting black hole stability, causal structure, and accretion dynamics [41, 40]. Recent research has explored magnetically charged black holes in NLE coupled to phantom fields, uncovering novel regular solutions and black bounce geometries [42]. Additionally, investigations into rotating and nonlinearly charged black holes with anisotropic matter fields have demonstrated significant deviations from standard GR solutions [43]. These modifications influence observable phenomena, including black hole shadows, gravitational lensing, and quasinormal modes (QNMs), providing potential avenues for testing deviations from classical GR [44]. Lower-dimensional studies, such as those on BTZ black holes, further highlight the unique thermodynamic and geometric properties introduced by NLE [39]. As research continues to explore the interplay between NLE and unconventional matter fields, these studies offer valuable insights into quantum gravity corrections, alternative compact objects, and broader modified gravity theories.
@@
Despite extensive research on phantom energy, nonlinear electrodynamics, and global monopoles individually, their combined impact on black hole physics remains largely unexplored. This study aims to fill this gap by systematically investigating how their interplay affects black hole thermodynamics, stability, and observational characteristics. Specifically, we construct and analyze a regular black hole solution sourced by a nonlinear magnetic field and pierced by a phantom global monopole. This setup reveals rich physical behavior, including nontrivial modifications to the horizon structure, thermodynamic phase transitions, and the spectrum of scalar quasinormal modes. Our results demonstrate that the combined effects of phantom topology and nonlinear electrodynamics introduce distinctive signatures that could, in principle, be probed observationally. To our knowledge, this is the first detailed investigation of such a composite system, highlighting its theoretical novelty and potential phenomenological relevance. This paper is structured as follows: Section 2 presents the derivation of the black hole solution and its geometric properties, while Section 3 analyzes its thermodynamic behavior. Section 4 investigates the timelike geodesics and the motion of test particles in the black hole spacetime. Section 5 examines QNMs using the sixth-order WKB approximation, cross-checked with the Pöschl-Teller potential. Finally, Section 6 analyzes grey-body factors and their influence on radiation emission, before we conclude in Section 7 with a summary and future outlook.
2 Spacetime Geometry and Black Hole Metric
@@
Following the approach of Barriola and Vilenkin [45], we assume that the scalar field function $h(r)$ quickly approaches unity outside the monopole core, since $h(r)$ grows linearly for $r<\delta$ and decays exponentially toward unity once $r>\delta$, where $\delta\sim(\eta\sqrt{\lambda})^{-1}$ defines the core radius of the global monopole. Under this assumption, the energy-momentum tensor associated with the global monopole takes the form:
@@
On the other hand, when Einstein gravity is coupled to an NLE field in the presence of a phantom global monopole, the total action can be written as:
$\mathcal{S}=\frac{1}{16\pi}\int\sqrt{-g}Rd^{4}x+\int d^{4}x\sqrt{-g}\Big(\mathcal{L}^{M}-\frac{1}{4\pi}\mathcal{L}^{charge}\Big),$ (6)
where the NLE Lagrangian density is given by:
Here, $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$, while $m$ and $Q$ correspond to the mass and magnetic charge of the system, respectively. Here, we introduce the magnetic charge $Q$ in the Lagrangian explicitly, since this mechanism is solution-dependent. For the present NLE case, this strategy is adopted and it follows reconstruction of the Lagrangian from a solution. A more general approach would be to introduce such a parameter and then relate it later; however, our choice is for analytic tractability and to elucidate the role played by the phantom monopole. In this work, we adopt a NLE Lagrangian that supports magnetically charged black hole solutions with desirable properties. The form given in Eq. (7) is inspired by regular black hole models such as the well-known Hayward solution, which has been shown to arise from specific NLE frameworks. Our aim is to investigate how this well-motivated NLE framework interacts with phantom global monopole fields, rather than to explore the most general possible NLE formulation. We assume that the scalar and NLE fields are minimally coupled, i.e., they interact only through the gravitational sector, and no direct coupling term is introduced between them. This choice is deliberate and well-motivated. The scalar field generates a global conical deformation in the asymptotic geometry, characteristic of global monopoles [45, 49], while the NLE field dominates the near-horizon behavior and is responsible for softening the central singularity [33, 35]. By keeping the two sectors minimally coupled, we ensure that their distinct effects remain analytically separable and theoretically transparent. Introducing a direct scalar-electromagnetic coupling would introduce new coupling parameters and complicate the interpretation of results, obscuring the independent roles of the scalar and electromagnetic sources.
Applying the variational principle from Eq. (6), the field equations governing the system take the form:
@@
The electromagnetic invariant remains strictly positive throughout the spacetime. This ensures that despite the non-integer exponents appearing in Eq. (7), the Lagrangian maintains both reality and regularity everywhere outside the origin. We note, however, that this form of the Lagrangian may become ill-defined in more general scenarios where $F$ could change sign, and this is a limitation of the specific ansatz we consider. Using the previously derived equations, the time component of Eq. (8) simplifies to:
$\frac{2}{r^{2}}\frac{dM\left(r\right)}{dr}=-\frac{6mQ^{3}}{\left(r^{3}+Q^{3}\right)^{2}}+8\pi\frac{\xi\eta^{2}}{r^{2}}.$ (14)
Integrating this expression, we obtain the metric function that describes a black hole with a NLE field in the presence of a phantom global monopole:
@@
We note that in the absence of the scalar field, the solution in Eq. (15) reduces precisely to the Hayward regular black hole [65]. Later the Hayward black hole was described by NLE theory in [66]. In the special case where $\xi\rightarrow 0$, this solution reduces to a black hole with a nonlinear magnetic charge. On the other hand, setting $Q\rightarrow 0$ retrieves the Schwarzschild black hole solution with a phantom global monopole [49]. We note that our choice of a regular NLE Lagrangian guarantees well-behaved field invariants and stress–energy if one extends the model toward small radii or matches to an interior solution. However, for the exterior analysis ($r\geq r_{c}$) presented here, the qualitative results remain unchanged for other standard NLE models (e.g., Maxwell, Born–Infeld, logarithmic), with only minor quantitative differences. This confirms that our main conclusions are robust against the specific NLE choice. It is crucial to highlight that NLE models, such as the square-root model—have been shown to induce angular deficits or excesses without invoking scalar fields [67]; however, these effects arise purely from the electromagnetic sector. In contrast, our work investigates the combined influence of two physically distinct sources: a phantom global monopole, which acts as a topological defect and modifies the large-scale structure of spacetime, and a nonlinear magnetic field, which governs the near-horizon geometry. These fields are minimally coupled, interacting only through gravity, which allows us to isolate their individual contributions while studying their collective impact on black hole physics. Although the inclusion of a phantom monopole renders the spacetime non–asymptotically flat, in this work we restrict attention to the exterior region $r\geq r_{c}$ relevant to the global–monopole setup. This setup provides a novel framework for exploring the interplay between topological charge and regularization mechanisms, with potential implications for black hole thermodynamics, quasinormal mode spectra, and gravitational wave phenomenology.
To investigate the curvature behavior of this spacetime, we compute the Ricci scalar $\mathbf{R}$ and the Kretschmann scalar $\mathbf{K}$, given by:
@@
As Eq. (17) shows, the Kretschmann scalar diverges as $r\to 0$, which means that the spacetime is not globally regular at the origin. In our analysis we only consider the exterior region $r\geq r_{c}$, where the geometry remains regular.
In Table 1 we include three NLE models that generate regular black holes, namely the Hayward, Bardeen, and Ayon-Beato-Garcia (ABG) spacetimes. Their metric functions, together with the corresponding curvature properties, are summarized there. We find that for $\eta=0$ all these models are regular at the origin, whereas for $\eta\neq 0$ the pointlike monopole contribution dominates and inevitably leads to a curvature divergence.
NED model Metric function $\lim\limits_{r\rightarrow 0}$ $\mathbf{K}$ ($\eta=0$) $\lim\limits_{r\rightarrow 0}$ $\mathbf{K}$ ($\eta\neq 0$)
Hayward $1-8\pi\xi\eta^{2}-\frac{2mr^{2}}{r^{3}+Q^{3}}$ $\frac{96m^{2}}{Q^{6}}$ $\infty$
Bardeen $1-8\pi\xi\eta^{2}-\frac{2mr^{2}}{\left(r^{2}+Q^{2}\right)^{3/2}}$ $\frac{96m^{2}}{Q^{2}}$ $\infty$
ABG $1-8\pi\xi\eta^{2}-\frac{2mr^{2}}{\left(r^{2}+Q^{2}\right)^{3/2}}+\frac{Q^{2}r^{2}}{\left(r^{2}+Q^{2}\right)^{2}}$ $\frac{24\left(2m-Q\right)^{2}}{Q^{6}}$ $\infty$
Table 1: Comparison of Hayward, Bardeen, and ABG BHs: metric functions and Kretschmann scalar limits with and without monopole.
@@
Moreover, the presence of the parameter $\xi$ (implicitly affecting $\eta$) influences the global behavior of the solution, indicating that a sufficiently strong monopole field can significantly alter black hole thermodynamics and causal structure. In particular, the limiting behavior at large $r$ remains consistent, although the spacetime is not asymptotically flat due to the solid-angle (conical) deficit induced by the monopole. However, near the black hole horizon, the modifications become more pronounced, potentially impacting observational signatures such as gravitational lensing and black hole shadows.
In this work, we adopt a specific form of NLE that reproduces the regular black hole solution proposed by Hayward [65] and later derived from a magnetic NLE source [66]. This choice ensures central regularity and allows for analytical control. We set the Schwarzschild mass parameter to zero, following the original Hayward formulation, where the effective mass emerges dynamically from the nonlinear magnetic sector. This assumption is particularly useful in our setup, as it allows us to isolate and analyze the gravitational interplay between the phantom global monopole and the NLE sector. While the phantom field alters the asymptotic structure through a topological angular deficit, the NLE Lagrangian governs the near-horizon geometry and ensures regularity at the core. The resulting spacetime, although not globally regular or asymptotically flat, offers a controlled framework to explore the combined geometric and thermodynamic effects of these two physically distinct sources.
@@ 3 Thermodynamics
where $\kappa$ represents the surface gravity at the event horizon. Although a global monopole introduces a solid-angle deficit, the Hawking temperature defined via the surface gravity, $T=\kappa/(2\pi)$ with $\kappa=\tfrac{1}{2}f^{\prime}(r_{H})$, is a local quantity at the horizon and remains valid. The angular deficit affects the horizon area (and thus the entropy), not the surface gravity.
To evaluate this temperature, we first determine the black hole mass $m$ in terms of the horizon radius $r_{H}$ by solving the condition $f(r_{H})=0$. This leads to the following expression:
@@
$T=\frac{\left(1-8\pi\xi\eta^{2}\right)}{4\pi r_{H}}\frac{\left(r_{H}^{3}-2Q^{3}\right)}{\left(r_{H}^{3}+Q^{3}\right)}.$ (20)
For the Hawking temperature to remain both real and positive, two conditions must be satisfied. The first constraint stems from the NLE field, establishing a lower bound on the event horizon radius $r_{H}$ in terms of the magnetic charge $Q$:
@@
Figure 2: Hawking temperature $T$ as a function of the horizon radius $r_{H}$ for different values of the symmetry-breaking scale $\eta$. The black hole parameters are set to $Q=0.5$ and $m=1$. The effect of the phantom global monopole ($\xi=-1$) leads to an increase in temperature, whereas the ordinary global monopole ($\xi=+1$) lowers it.
@@
where $A=4\pi{(1-8\pi\xi\eta^{2})}r_{H}^{2}$ is the physical horizon area rescaled by the solid–angle deficit due to a global monopole. It is noteworthy that the entropy does not exactly coincide with the standard Bekenstein–Hawking law. In particular, Eq.  (24) contains an additional correction term arising from NLE. The standard Bekenstein–Hawking form is recovered either when $\xi\!\to\!0$ and $Q\!\to\!0$ (the conventional case without a monopole and without NLE), or, approximately, in the large–area limit where the NLE correction becomes negligible.
In addition to the temperature and entropy analysis, we derive a first law and a Smarr-like formula for the nonlinearly magnetically charged black hole in the presence of a phantom global monopole. From Eq. (19), we express the black hole mass $m$ in terms of the horizon radius $r_{H}$, magnetic charge $Q$ and symmetry-breaking scale $\eta$. Differentiating $m$, we find
$dm=TdS+\Psi dQ+\Phi d\eta,$ (25)
where
$\Psi=\left.\frac{\partial m}{\partial Q}\right|_{r_{H},\eta}=\frac{3r_{H}}{2}\left(1-8\pi\xi\eta^{2}\right)\frac{Q^{2}}{r_{H}^{3}},$ (26)
and
$\Phi=\left.\frac{\partial m}{\partial{\eta}}\right|_{r_{H},Q}=-r_{H}\left(1+\frac{Q^{3}}{r_{H}^{3}}\right)8\pi\xi\eta,$ (27)
represent the generalized potentials conjugate to $Q$ and $\eta$, respectively. These terms account for the contributions of the NLE and the phantom global monopole to the black hole’s energy content. Furthermore, using Euler’s theorem for homogeneous functions, we obtain a Smarr-type formula:
$m=2TS+\Psi Q-\Phi\eta.$ (28)
@@
$G=\frac{\left(1-8\pi\eta^{2}\xi\right)\left(r_{H}^{6}+8Q^{3}r_{H}^{3}-2Q^{6}\right)}{4\,r_{H}^{2}\left(r_{H}^{3}+Q^{3}\right)}.$ (32)